# POD-Attention: Unlocking Full Prefill-Decode Overlap for Faster LLM Inference

### v1

@@ Abstract.
Each request in LLM inference goes through two phases: compute-bound prefill and memory-bandwidth-bound decode. To improve GPU utilization, recent systems use hybrid batching that combines the prefill and decode phases of different requests into the same batch. Hybrid batching works well for linear operations as it amortizes the cost of loading model weights from HBM. However, attention computation in hybrid batches remains inefficient because existing attention kernels are optimized for either prefill or decode.
In this paper, we present POD-Attention [^1] Available at https://github.com/microsoft/vattention/tree/main/pod_attn. — the first GPU kernel that efficiently computes attention for hybrid batches. POD-Attention aims to maximize the utilization of both compute and memory bandwidth by carefully allocating the GPU’s resources such that prefill and decode operations happen concurrently on the same multiprocessor. We integrate POD-Attention in a state-of-the-art LLM inference scheduler Sarathi-Serve. POD-Attention speeds up attention computation by up to $75\%$ (mean 28%) and increases LLM serving throughput by up to $22\%$ in offline inference. In online inference, POD-Attention enables lower time-to-first-token (TTFT), time-between-tokens (TBT), and request execution latency versus Sarathi-Serve.
@@ 1. Introduction
The serving infrastructure of large language models (LLMs) is expanding to meet growing demands (ai-infra-spending, ; ai-infra-spending-2, ). Large-scale service providers often depend on expensive high-end GPUs to meet peak demand or latency targets (patel2023splitwise, ). Therefore, optimizing LLM serving systems has become crucial (orca, ; vllmsosp, ; sarathiserve2024, ; distserve2024, ; nanoflow2024, ; loongserve2024, ; mnemosyne2024, ). Importantly, the overall efficiency of a deployment depends on how well GPU resources are utilized.
From a resource utilization perspective, LLM inference is a challenging workload because different phases require different resources at different times (sarathi2023, ; sarathiserve2024, ; nanoflow2024, ; vidur, ). The processing of an LLM request begins with a highly parallel (hence compute-bound) prefill phase which is then followed by a memory-bound decode phase  (sarathi2023, ). Serving LLMs efficiently, therefore, requires both high compute and high memory bandwidth. An ideal system would strive to maximize the utilization of both compute and memory. However, doing so is non-trivial because, for a given request, the prefill and decode phases occur at different times.
@@
State-of-the-art LLM serving systems deal with this challenge by combining the inputs of prefill and decode phases of different requests into the same batch (sarathi2023, ; splitfuse2024, ; orca, ) — a technique we refer to as hybrid batching. Hybrid batching avoids the need to fetch model weights from GPU high-bandwidth memory (HBM) separately for prefill and decode tokens. Instead, it allows the GPU to fetch model weights once and use them to compute over both prefill and decode inputs. Hybrid batching also helps reduce tail latency: to limit the runtime of each iteration, the scheduler can divide long input prompts (prefill inputs) into multiple smaller chunks, then combine ongoing decodes with a new prefill chunk every iteration (sarathiserve2024, ; splitfuse2024, ). As such, use of hybrid batching is common in various LLM serving systems today (vllmsosp, ; nanoflow2024, ; sarathiserve2024, ; orca, ; splitfuse2024, ).
@@
While prior work has focused on optimizing the linear operations (sarathiserve2024, ; orca, ; splitfuse2024, ), they do not optimize the attention computation of a hybrid batch. This is reasonable for a system that primarily deals with small context lengths since linear operations dominate run time in this setting (orca, ; nanoflow2024, ). In contrast, as the context length increases, attention computation becomes the primary performance bottleneck (Figure 4).
@@
Some recent works have also tried to optimize attention computation (flashattention, ; flashdecoding, ; flashdecoding++, ; leanattention, ), but current solutions address prefill and decode operations separately — maximizing compute utilization for prefills and bandwidth utilization for decodes, as shown in Figure 1. In this paper, we show that such an approach is suboptimal as it leaves critical GPU resources underutilized in different parts of computation. For example, Figure 1 illustrates that memory bandwidth utilization of the prefill attention kernel is often below 5%, while compute utilization of the decode attention kernel is under 10%. The effect of using independently optimized kernels is particularly noticeable with hybrid batching because prefill and decode kernels execute immediately one after the other, leading to periods of high demand of a resource immediately followed by low utilization of the same resource.
@@
To improve the efficiency of hybrid batching, we present POD-Attention — the first GPU kernel, to the best of our knowledge, that efficiently batches the computation of prefill and decode attention. In doing so, we first show (§3) that existing techniques do not provide adequate performance in fusing attention computation due to various limitations such as straggler threads, synchronization barriers and lack of guaranteed SM-level co-location of different Cooperative Thread Arrays (CTAs) on GPU Streaming Multiprocessors (SMs). POD-Attention addresses these issues by fusing the computation in a CTA-parallel manner, introducing SM-aware software-based CTA scheduling within the GPU (§4). Building on state-of-the-art FlashAttention kernels (fagithub, ), POD-Attention significantly accelerates attention computation by utilizing both compute and memory resources as per the requirement of a given batch of requests (see Figure 1).
@@
We also integrate POD-Attention in a state-of-the-art LLM inference scheduler Sarathi-Serve (sarathiserve2024, ). Our experiments show that POD-Attention computes attention up to $75\%$ faster (mean 28%) than the prefill and decode attention kernels of FlashAttention and FlashInfer. In terms of the end-to-end LLM inference performance, POD-Attention improves throughput by up to $22\%$ while also reducing crucial latency metrics such as time-to-first-token (TTFT), time-between-tokens (TBT) and the end-to-end request execution latency over Sarathi-Serve.
@@
  • We highlight that independently optimizing prefill and decode attention kernels is not performant with hybrid-batching-based LLM inference.
  • We present POD-Attention — a single GPU kernel that computes prefill and decode attention concurrently to try and utilize both compute and memory bandwidth simultaneously.
  • We integrate POD-Attention in Sarathi-Serve and show that it enables high throughput and low latency LLM inference compared to current systems.
@@ 2. Background and Motivation
2.1. Large Language Model (LLM) Inference
LLMs process user inputs and outputs as tokens, internally represented as vectors. Each request during inference goes through two phases — prefill and decode (orca, ). The prefill phase processes the tokens of a user’s prompt in parallel and produces the first output token, whose latency is called time-to-first-token (TTFT). Subsequently, the decode phase generates one output token (per-request) per-iteration auto-regressively. The latency taken to generate each output token is called time-between-tokens (TBT). The prefill phase is highly parallel and compute bound while the decode phase is memory bound. Due to the parallel processing of a large number of tokens, the latency of a prefill iteration is generally higher than that of a decode iteration.
The distinct computational characteristics of prefill and decode operations create a throughput-latency tradeoff in LLM inference (sarathiserve2024, ; patel2023splitwise, ; distserve2024, ; tetriinfer, ), as illustrated in Figure 2. Since decoding is memory bound, using a large batch size improves throughput. The original vLLM scheduler (vllmsosp, ) uses prefill-prioritizing scheduling to maximize the decode batch size (Figure 2(a)). This approach provides low TTFT, but at the cost of high TBT because a new request’s prefill can pause ongoing decodes, causing generation stalls (sarathiserve2024, ). High TBT is especially problematic in long-context scenarios, where each generation stall can last several seconds.
The issue of high TBT has been acknowledged in real-world deployments (vllm:performance-guide-chunking, ). Sarathi-Serve (sarathiserve2024, ) proposed chunked-prefills coupled with continuous hybrid batching (orca, ) — a technique that divides the prefill tokens of a request into multiple smaller chunks and schedules one prefill chunk per-iteration with on-going decodes (Figure 2(b)). This way, Sarathi-Serve enables increasing batch size while avoiding generation stalls, improving both performance and user interactivity. Various LLM serving systems have incorporated this technique (nanoflow2024, ; sglang2024, ; trtllmgithub, ) including vLLM (vllm:upstream-chunked-prefill, ). However, a downside is that it increases TTFT because each prefill chunk experiences interference from co-running decodes (sarathi2023, ). Since longer contexts are split into more chunks, the TTFT of long context requests can become significantly higher.
Figure 3 shows how hybrid batching works. Except attention, all other operations are linear i.e., computed element-wise. Linear operations obey the rule f(x + y) = f(x) + f(y) so inputs for a linear operation can be combined, computed upon by the same model weights to reduce memory accesses, and then separated. In contrast, attention is a sequence-level operator that is computed between three representations Q (query, of the current tokens being processed), and K/V (key/value, of all tokens in the sequence seen so far) as:
@@ 2.1. Large Language Model (LLM) Inference
The QKV representations are further divided among multiple query heads and K/V heads, each assigned to a group (ainslie2023gqa, ). Attention is computed in parallel for each Q head and K/V head pair. Since resource requirements of prefill and decode attention are different, state-of-the-art libraries such as FlashAttention (FA) (flashattention, ; flashattention2, ; flash-attention-3, ) and FlashInfer (FI) (flashinfer, ) provide specialized kernel APIs, optimized separately for each phase. Use of these kernels works well in small context length scenarios where attention computation is a small fraction of the total inference time (sarathi2023, ; orca, ).
However, the context length in many real-world LLM applications continues to grow (mnemosyne2024, ; loongserve2024, ). In such scenarios, attention computation dominates, becoming more than 60% of the total inference time in many cases as shown in Figure 4 (context length 16K). Note that prefill and decode attention are computed immediately one after the other in hybrid batches (see Figure 3). Therefore, when independently optimized attention kernels are used, GPU execution goes through periods of high demand of a resource followed by low utilization of the same resource. For example, the prefill kernel requires high compute but compute is (mostly) idle when the decode kernel executes.
@@ 2.2. GPU Execution Model
The GPU’s hardware is arranged in a hierarchy that supports execution at a scale of hundreds of thousands of parallel threads, depicted in Figure 5 (CUDA_HW, ). The main processor unit of a GPU is a Streaming Multiprocessor (SM), with modern GPUs containing around a hundred SMs. Each SM has an L1 cache and shared memory along with tensor cores for accelerated general matrix multiplication (GEMM) and execution units for integer/floating point operations. The shared memory is a user-addressable partition of the L1 cache. The GPU memory is accessed by SMs through the shared L2 cache.
GPU programming languages expose a hierarchy of threads that mimic the hardware hierarchy. The smallest unit of execution is a thread, while a group of 32 threads make up a warp, which typically execute concurrently in lockstep. To maximize throughput, GPU programmers ensure that threads within a warp execute the same code path. A Cooperative Thread Array (CTA) (PTX_CTA, ) is a group of warps that share the L1 cache and shared memory. All warps in a CTA are guaranteed to execute within a single SM.
@@ 3. A Case Study on Concurrent Execution
The simplest way to compute prefill and decode attention together is to pass both inputs to an existing attention kernel. Some LLM serving systems prefer this method for computing attention in hybrid batches (vllm:unifyattentionkernel, ; sarathi:unifiedattentionkernel, ). In §5.1, we show that this is counter-productive and slower than serial execution.
@@ 3.1. Methods of Concurrent Execution
    Kernel-parallel. Streams can potentially execute different GPU kernels concurrently. This approach is easy to implement as it only requires submitting existing kernels to different streams; all other approaches require fusing different operations into a single kernel. Unfortunately, streams alone guarantees neither concurrency nor SM-level co-location of different operations (Elastic_kernel_ASPLOS13, ; ISPA_TOC23, ).
@@
    Streams (Elastic_kernel_ASPLOS13, ) $\times$ Easiest to implement
@@
    Warp (e.g., HFuse (HFuse_CGO_2022, )) $\times$ Suffers from straggler problem
    Intra-thread (kernel_fuse, ; kernel_waver, ) $\times$ Cannot overlap with CTA barriers
@@
    FA_HFuse Horizontally fused FA kernels with HFuse (HFuse_CGO_2022, )
@@
    Warp-parallel. Here, warps within each CTA are split across operations, as proposed in horizontal fusion (HFuse (HFuse_CGO_2022, )). This apprach guarantees co-location since all warps in a CTA are guaranteed to reside within the same SM. Unfortunately, warp-parallel fusion suffers from the straggler problem: a CTA prevents other CTAs from being scheduled on the same SM if one or more of its threads/warps are lagging behind others.
@@
    Intra-thread. In intra-thread fusion, each thread alternates between executing instructions of different operations (kernel_fuse, ; kernel_waver, ). In simple cases, this strategy provides the maximum opportunity to overlap different operations. However, attention kernels use CTA-level sync barriers to coordinate fetching data into shared memory. These barriers limit intra-thread fusion as instructions before a barrier cannot be overlapped with those after the barrier.

### v2

@@ Abstract.
Each request in LLM inference goes through two phases: compute-bound prefill and memory-bandwidth-bound decode. To improve GPU utilization, recent systems use hybrid batching that combines the prefill and decode phases of different requests into the same batch. This approach optimizes linear operations but remains inefficient for attention computation because existing attention kernels specialize execution independently for the prefill and decode phases.
In this paper, we present POD-Attention — the first GPU kernel that efficiently computes attention for hybrid batches. POD-Attention aims to maximize the utilization of both compute and memory bandwidth by carefully allocating the GPU’s resources such that prefill and decode operations happen concurrently on the same multiprocessor. POD-Attention speeds up attention computation by up to $59\%$ (mean 28%), enabling higher throughput and lower latency LLM inference compared to the use of independently optimized prefill and decode attention kernels.
Large language models; GPUs; self-attention
journalyear: 2025copyright: ccconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 30-April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’25), March 30-April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3676641.3715996isbn: 979-8-4007-1079-7/2025/03ccs: Computing methodologies Machine learningccs: Computer systems organization Parallel architectures
@@ 1. Introduction
The infrastructure for serving large language models (LLMs) is expanding to meet their growing demands (ai-, 2024a, b). Large-scale service providers often depend on expensive high-end GPUs to meet peak demand or latency targets (Patel et al., 2024). Therefore, optimizing LLM serving systems has become crucial (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024c; Zhong et al., 2024; Zhu et al., 2024; Wu et al., 2024; Agrawal et al., 2024a). The overall efficiency of a deployment depends on how well GPU resources are utilized.
From a resource utilization perspective, LLM inference is a challenging workload because different phases require different resources at different times (Agrawal et al., 2023, 2024c; Zhu et al., 2024; Agrawal et al., 2024b). The processing of an LLM request begins with a highly parallel (hence, compute-bound) prefill phase which is then followed by a memory-bound decode phase  (Agrawal et al., 2023). Serving LLMs efficiently, therefore, requires both high compute and high memory bandwidth. An ideal system would strive to maximize the utilization of both compute and memory. However, doing so is non-trivial because for a given request, the prefill and decode phases occur at different times.
@@
State-of-the-art LLM serving systems deal with this challenge by combining the inputs of prefill and decode phases of different requests into the same batch (Agrawal et al., 2023; Holmes et al., 2024; Yu et al., 2022) — a technique we refer to as hybrid batching. Hybrid batching avoids the need to fetch model weights from GPU high-bandwidth memory (HBM) separately for prefill and decode tokens. Instead, it allows the GPU to fetch model weights once and use them to compute over both prefill and decode inputs. Hybrid batching also helps reduce tail latency: to limit the runtime of each iteration, the scheduler can divide long input prompts (prefill inputs) into multiple smaller chunks, then combine ongoing decodes with a new prefill chunk every iteration (Agrawal et al., 2024c; Holmes et al., 2024). As such, use of hybrid batching is common in various LLM serving systems today (Kwon et al., 2023; Zhu et al., 2024; Agrawal et al., 2024c; Yu et al., 2022; Holmes et al., 2024).
@@
While prior work has focused on optimizing the linear operations (Agrawal et al., 2024c; Yu et al., 2022; Holmes et al., 2024), they do not optimize the attention computation of a hybrid batch. This is reasonable for a system that primarily deals with small context lengths since linear operations dominate run time in this setting (Yu et al., 2022; Zhu et al., 2024). In contrast, as the context length increases, attention computation becomes the primary performance bottleneck (Figure 4).
@@
Some recent works have also tried to optimize attention computation (Dao et al., 2022, 2023; Hong et al., 2024; Sanovar et al., 2024), but current solutions address prefill and decode operations separately — maximizing compute utilization for prefills and bandwidth utilization for decodes, as shown in Figure 1. In this paper, we show that such an approach is suboptimal as it leaves critical GPU resources underutilized in different parts of computation. For example, Figure 1 illustrates that memory bandwidth utilization of the prefill attention kernel is often below 5%, while compute utilization of the decode attention kernel is under 10%. The effect of using independently optimized kernels is particularly noticeable with hybrid batching because prefill and decode kernels execute immediately one after the other, leading to periods of high demand of a resource immediately followed by low utilization of the same resource.
@@
To improve the efficiency of hybrid batching, we present POD-Attention — the first GPU kernel, to the best of our knowledge, that efficiently batches the computation of prefill and decode attention. In doing so, we first show (§3) that existing techniques do not provide adequate performance in fusing attention computation due to various limitations such as straggler threads, synchronization barriers and lack of guaranteed SM-level co-location of different Cooperative Thread Arrays (CTAs) on GPU Streaming Multiprocessors (SMs). POD-Attention addresses these issues by fusing the computation in a CTA-parallel manner, introducing SM-aware software-based CTA scheduling within the GPU (§4). Building on state-of-the-art FlashAttention kernels (fag, 2022), POD-Attention significantly accelerates attention computation by utilizing both compute and memory resources as per the requirement of a given batch of requests (see Figure 1).
@@
We also integrate POD-Attention in a state-of-the-art LLM inference scheduler Sarathi-Serve (Agrawal et al., 2024c). Our experiments show that POD-Attention computes attention up to $59\%$ faster (mean 28%) than the prefill and decode attention kernels of FlashAttention and FlashInfer. In terms of the end-to-end LLM inference performance, POD-Attention improves throughput by up to $22\%$ while also reducing crucial latency metrics such as time-to-first-token (TTFT), time-between-tokens (TBT) and the end-to-end request execution latency over Sarathi-Serve.
@@
  • We highlight that independently optimizing prefill and decode attention kernels is suboptimal for hybrid batching based LLM inference.
  • We present POD-Attention — a GPU kernel that computes prefill and decode attention concurrently to utilize both compute and memory bandwidth simultaneously.
  • We integrate POD-Attention in Sarathi-Serve and show that it enables high throughput and low latency LLM inference compared to the use of independently optimized prefill and decode attention kernels.
@@ 2. Background and Motivation
2.1. Large Language Model (LLM) Inference
LLMs process user inputs and outputs as tokens, internally represented as vectors. Each request during inference goes through two phases — prefill and decode (Yu et al., 2022). The prefill phase processes the tokens of a user’s prompt in parallel and produces the first output token, whose latency is called time-to-first-token (TTFT). Subsequently, the decode phase generates one output token (per-request) per-iteration auto-regressively. The latency taken to generate each output token is called time-between-tokens (TBT). The prefill phase is highly parallel and compute bound while the decode phase is memory bound. Due to the parallel processing of a large number of tokens, the latency of a prefill iteration is generally higher than that of a decode iteration.
@@ 2.1. Large Language Model (LLM) Inference
The distinct computational characteristics of prefill and decode operations create a throughput-latency tradeoff in LLM inference (Agrawal et al., 2024c; Patel et al., 2024; Zhong et al., 2024; Hu et al., 2024), as illustrated in Figure 2. Since decoding is memory bound, using a large batch size improves throughput. The original vLLM scheduler (Kwon et al., 2023) uses prefill-prioritizing scheduling to maximize the decode batch size (Figure 2(a)). This approach provides low TTFT, but at the cost of high TBT because a new request’s prefill can pause ongoing decodes, causing generation stalls (Agrawal et al., 2024c). High TBT is especially problematic in long-context scenarios, where each generation stall can last several seconds.
The issue of high TBT has been acknowledged in real-world deployments (vll, 2024a). Sarathi-Serve (Agrawal et al., 2024c) proposed chunked-prefills coupled with continuous hybrid batching (Yu et al., 2022) — a technique that divides the prefill tokens of a request into multiple smaller chunks and schedules one prefill chunk per-iteration with on-going decodes (Figure 2(b)). This way, Sarathi-Serve enables increasing batch size while avoiding generation stalls, improving both performance and user interactivity. Various LLM serving systems have incorporated this technique (Zhu et al., 2024; Zheng et al., 2024; trt, 2023), including vLLM (vll, 2024c).
In the common case with hybrid batching, an executing batch consists of one prefill chunk of a pre-determined size and multiple decodes (as shown in Table 1). For example, consider a workload where each request consists of 2K prefill tokens and generates 200 output (decode) tokens. If the prefill chunk size is 1K, a request’s prefill completes over two iterations (prefill tokens / chunk size). Upon completion of the prefill phase, it must execute for another 200 iterations — each iteration corresponding to one output token. In these 200 iterations, 100 requests can complete their prefill phase to join the running batch. This leads to an effective batch size of 101 in the steady state wherein 100 requests execute in their decode phase alongside one prefill chunk of a new request. Executing these hybrid batches requires both high compute (for the prefill chunk) and high memory bandwidth (for the decode requests).
Figure 3 shows how hybrid batching works in practice. Except attention, all other operations are linear i.e., computed element-wise. Linear operations obey the rule f(x + y) = f(x) + f(y) so inputs for a linear operation can be combined, computed upon by the same model weights to reduce memory accesses, and then separated. In contrast, attention is a sequence-level operator that is computed between three representations Q (query, of the current tokens being processed), and K/V (key/value, of all tokens in the sequence seen so far) as:
@@
The QKV representations are further divided among multiple query heads and K/V heads, each assigned to a group (Ainslie et al., 2023). Attention is computed in parallel for each Q head and K/V head pair. Since resource requirements of prefill and decode attention are different, state-of-the-art libraries such as FlashAttention (FA) (Dao et al., 2022; Dao, 2024; Shah et al., 2024) and FlashInfer (FI) (Ye et al., 2025) provide specialized kernel APIs, optimized separately for each phase. Use of these kernels works well in small context length scenarios where attention computation is a small fraction of the total inference time (Agrawal et al., 2023; Yu et al., 2022).
However, the context length in many real-world LLM applications continues to grow (Agrawal et al., 2024a; Wu et al., 2024). In such scenarios, attention computation dominates, becoming more than 60% of the total inference time in many cases as shown in Figure 4 (context length 16K). Note that prefill and decode attention are computed immediately one after the other in hybrid batches (see Figure 3). Therefore, when independently optimized attention kernels are used, GPU execution goes through periods of high demand of a resource followed by low utilization of the same resource. For example, the prefill kernel requires high compute but compute is (mostly) idle when the decode kernel executes.
@@ 2.2. GPU Execution Model
The GPU’s hardware is arranged in a hierarchy that supports execution at a scale of hundreds of thousands of parallel threads, depicted in Figure 5 (CUD, 2024). The main processor unit of a GPU is a Streaming Multiprocessor (SM), with modern GPUs containing around a hundred SMs. Each SM has an L1 cache and shared memory along with tensor cores for accelerated general matrix multiplication (GEMM) and execution units for integer/floating point operations. The shared memory is a user-addressable partition of the L1 cache. The GPU memory is accessed by SMs through the shared L2 cache.
GPU programming languages expose a hierarchy of threads that mimic the hardware hierarchy. The smallest unit of execution is a thread, while a group of 32 threads make up a warp, which typically execute concurrently in lockstep. To maximize throughput, GPU programmers ensure that threads within a warp execute the same code path. A Cooperative Thread Array (CTA) (PTX, 2024a) is a group of warps that share the L1 cache and shared memory. All warps in a CTA are guaranteed to execute within a single SM.
@@ 3. A Case Study on Concurrent Execution
The simplest way to compute prefill and decode attention together is to pass both inputs to an existing attention kernel. Some LLM serving systems prefer this method for computing attention in hybrid batches (vll, 2024b; sar, 2024a). In §5.1, we show that this is counter-productive and slower than serial execution.
@@ 3.1. Methods of Concurrent Execution
    Kernel-parallel. Streams can potentially execute different GPU kernels concurrently. This approach is easy to implement as it only requires submitting existing kernels to different streams; all other approaches require fusing different operations into a single kernel. Unfortunately, streams alone guarantees neither concurrency nor SM-level co-location of different operations (Pai et al., 2013; Zhao et al., 2023).
@@
    Streams (Pai et al., 2013) $\times$ Easiest to implement
@@
    Warp (e.g., HFuse (Li et al., 2022)) $\times$ Suffers from straggler problem
    Intra-thread (Wahib and Maruyama, 2014; Wu et al., 2012) $\times$ Cannot overlap with CTA barriers
@@
    FA_HFuse Horizontally fused FA kernels with HFuse (Li et al., 2022)
@@
    Warp-parallel. Here, warps within each CTA are split across operations, as proposed in horizontal fusion (HFuse (Li et al., 2022)). This apprach guarantees co-location since all warps in a CTA are guaranteed to reside within the same SM. Unfortunately, warp-parallel fusion suffers from the straggler problem: an entire CTA must complete execution before it can be replaced by another one; if one or more of its threads or warps lag behind others, the next CTA is delayed. While fusing the prefill and decode attention computation, the fused kernel requires extensive tuning to deal with a large input space of varying batch sizes and context lengths e.g., some hybrid batches may be prefill heavy and others may be decode heavy. Therefore, a fused prefill-decode attention kernel is particularly vulnerable to the straggler effect with warp-parallel fusion.
@@
    Intra-thread. In intra-thread fusion, each thread alternates between executing instructions of different operations (Wahib and Maruyama, 2014; Wu et al., 2012). In simple cases, this strategy provides the maximum opportunity to overlap different operations. However, attention kernels use CTA-level sync barriers to coordinate fetching data into shared memory. These barriers limit intra-thread fusion as instructions before a barrier cannot be overlapped with those after the barrier.